Find three consecutive odd integers such that two times the sume of the first and the third integer is thirteen more than the product of three and the second number.
I did it like this, help??????
2x(x+x+4)=13+(3+x+2)
12x=help

Answer by Paul(988)

First = x
Second = x+2
Third = x+4
2(x+x+4)=13+3(x+2)
4x+8=13+3x+6
4x+8=3x+19
x=11
11+2=13
11+4=15
Hence, the three consective odds integers are 11, 13, and 15.